OpenStudy (anonymous):

Determine the domain, range, zeros, y intercepts, end behavior, and intervals of increasing or decreasing of f(x) = x(x+4)(x+5)^3

OpenStudy (campbell_st):

well do you have any thoughts... about the question... or have some answers...?

OpenStudy (anonymous):

i have the domain, range, and end behavior, but i don't know how to get the other parts

OpenStudy (campbell_st):

ok... so the zeros... the polynomial is factored so you need to solve x = 0 x+4 = 0 x +5 = 0 thats a start

OpenStudy (anonymous):

0, -4, -5, thanks! i have no idea what to do for y intercepts and I'm trying increasing/decreasing on my calculator but I'm still kinda confused on it

OpenStudy (campbell_st):

ok... so the y intercept is 0... since the product of the constants will give the y intercept so the polynomial can be written as \[y = (x + 0)(x +4)(x + 5)^3\] the constants are 0, 4 and 5^3 so their product is \[0 \times 4 \times 5^3= ?\] what do you think

OpenStudy (anonymous):

that would equal 0

OpenStudy (campbell_st):

yep... so thats the y-intercept..

OpenStudy (anonymous):

thank you!! you're really smart and i appreciate how you are making me think and not just giving me the answer, it's really helpful

OpenStudy (campbell_st):

ok... so the function is an odd power... starting lower left and finishing upper right... so it looks something like this....  you'll probably need to use calculus to find the 1st derivatives.... and max and min values... and 2nd derivatives to find the points of inflexion